# appendixapcList of Appendices

### v2

@@ Abstract
Large language models (LLMs) have shown remarkable capabilities, but still struggle with processing extensive contexts, limiting their ability to maintain coherence and accuracy over long sequences. In contrast, the human brain excels at organising and retrieving episodic experiences across vast temporal scales, spanning a lifetime. In this work, we introduce EM-LLM, a novel approach that integrates key aspects of human episodic memory and event cognition into LLMs with no fine-tuning, enabling them to handle practically infinite context lengths while maintaining computational efficiency. EM-LLM organises sequences of tokens into coherent episodic events using a combination of Bayesian surprise and graph-theoretic boundary refinement in an online fashion. When needed, these events are retrieved through a two-stage memory process, combining similarity-based and temporally contiguous retrieval for efficient and human-like access to relevant information. Experiments on the LongBench and $\infty$-Bench benchmarks demonstrate EM-LLM’s superior performance, consistently outperforming the state-of-the-art retrieval model InfLLM across various baseline LLMs. In addition, EM-LLM outperforms its popular counterpart, RAG, in a wide range of tasks, while requiring similar resources. Notably, EM-LLM’s performance even surpasses full-context models in most tasks, while successfully performing retrieval across 10 million tokens -- a scale computationally infeasible for such models. Finally, our analysis reveals strong correlations between EM-LLM’s event segmentation and human-perceived events, suggesting a bridge between this artificial system and its biological counterpart, thereby offering a novel computational framework for exploring human memory mechanisms.
@@ 1 Introduction
Figure 1: Top: EM-LLMS vs. RAG (NV-Embed-v2 retriever) vs. full-context, with LLaMA-3.1-8B as the base LLM, evaluated on LongBench. Bottom: Comparison of various long-sequence methods (sorted based on their context window length) on an extended version of $\infty$-Bench’s Retrieve.PassKey. Baseline data taken from Ding et al. [2024].
@@
Following these insights, we propose a novel architecture, EM-LLM, which integrates crucial aspects of event cognition and episodic memory into Transformer-based LLMs. For memory formation, we segment the sequence of tokens presented to the underlying LLM into individual memory units representing episodic events. The boundaries, and thus the size of those events, are initially determined dynamically, based on the level of surprise of the model during inference, and then refined to maximise cohesion within memory units and separation of memory content across them (see Section 3.2). This refinement process leverages graph-theoretic metrics, treating the similarity between attention keys (the learned representations used in Transformer self-attention mechanisms) as a weighted adjacency matrix, and aims to enhance the model’s ability to efficiently recall relevant information when addressing complex tasks with extended contexts. Importantly, this memory formation process incurs minimal additional computational cost. The surprise-based segmentation requires no extra computation and the refinement step has a complexity of $\mathcal{O}(kn)$, where $k$ is typically very small compared to the number of tokens $n$. With this efficient memory formation process, by grouping similar information in single units, we minimise the number of units needed to recall details around specific events. For memory recall, our approach integrates similarity-based retrieval with mechanisms that facilitate temporal contiguity and asymmetry effects. By retrieving and buffering salient memory units, our model leverages and enhances the recently discovered propensity of LLMs to exhibit human-like patterns in sequential information retrieval [Ji-An et al., 2024]. This method not only ensures efficient access to pertinent information but also mimics the temporal dynamics found in human free recall studies (such as Howard and Kahana [2002]), further enhancing the model’s ability to handle complex tasks that require nuanced temporal reasoning.
@@
In this work, we show that our method is scalable and significantly outperforms the SOTA retrieval model InfLLM [Xiao et al., 2024a], as well as RAG and full-context methods, on the widely-used LongBench [Bai et al., 2023] and $\infty$-Bench [Zhang et al., 2024] benchmarks designed for long-context tasks (see Fig. 1). Furthermore, we perform successful passkey retrieval across $10$M tokens, a length which is computationally infeasible for current full-context models. To further prove our hypotheses, we then employ a series of human-annotated podcast scripts to show that information in LLM attention heads can be semantically grouped in a way that correlates with the event structure perceived by humans. Therefore, LLM-perceived surprise can indeed serve as a proxy for the cognitive signals that drive human event segmentation, as confirmed by previous studies [Kumar et al., 2023]. Finally, using the long-context PG-19 dataset [Rae et al., 2020], which comprises a diverse corpus of English books, we evaluate the effectiveness of our segmentation method for grouping relevant information and assess the performance of different boundary refinement objectives.
@@ 2.1 Long-context in LLMs
Recently, several approaches have been proposed to extend the context window of Transformer-based models. These include methods that address the limited representational capacity of softmax attention, and its quadratic computational and memory cost [Katharopoulos et al., 2020, Munkhdalai et al., 2024]. Other methods target the poor extrapolation of typical positional encodings to out-of-distribution context lengths [Kazemnejad et al., 2024]. The latter is evident in most widely used methods, including the original absolute positional encodings [Vaswani et al., 2017] and the more recent relative positional encodings, such as the popular Rotary Positional Embeddings (RoPE) [Su et al., 2024]. To address this, some approaches propose scaling of the rotation angles [Chen et al., 2024a] or the base constant [Xiong et al., 2023, Liu et al., 2024b, Peng et al., 2024, Ding et al., 2024]. Others, scale positions without affecting the embedding function [Press et al., 2021, Chen et al., 2023, Jin et al., 2024], explore alternative strategies such as KERPLE [Chi et al., 2022] and FIRE [Li et al., 2024a] or adopt relative position embedding mechanisms from certain LMs like T5 [Raffel et al., 2020].
@@
Concerning computational efficiency and diluted attention, successful approaches propose methods for general improvements to Transformer efficiency through optimised computations [Dao, 2024, Han et al., 2024a, Aminabadi et al., 2022, Kwon et al., 2023, Liu et al., 2024c, Brandon et al., 2023] or compression techniques [Nawrot et al., 2024, Zhang et al., 2023], as well as training methods tailored for long-context scenarios [Zhu et al., 2024, Chen et al., 2024b]. Another direction is the utilisation of retrieval-based methods, the vast majority of which relies on a vector database that keeps a key-value cache and scalable approximations of k-nearest neighbors (k-NNs) to perform lookups [Wu et al., 2022, Tworkowski et al., 2023, Bertsch et al., 2023]. Interestingly, since using a key-value cache with k-NN lookup can be seen as an approximation of applying softmax attention to the full token sequence (see Appendix E.1), k-NN retrieval methods can be used without fine-tuning [Bertsch et al., 2023]. For an exception that does not rely on k-NNs, see Wang et al. [2023].
@@
A recent and interesting variant of k-NN retrieval involves retrieving large groups of tokens, rather than individual ones. Models that rely on this approach include SLED [Ivgi et al., 2023] and the more recent InfLLM [Xiao et al., 2024a], which achieves SOTA performance on long-context benchmarks. InfLLM segments the entire context length into fixed-size memory units and employs k-NN lookup using the tokens with the highest accumulated scores per unit. This can be seen as a form of hierarchical attention, as illustrated in Fig. 2. While group-based retrieval represents a promising direction, our approach significantly advances this concept by dynamically determining token groupings in a manner akin to human memory formation. This effectively addresses a fundamental limitation of InfLLM’s fixed-size segmentation and enables more adaptive and context-sensitive processing of extended information.
@@ 3.2 Memory formation via Surprise
$-\log P(x_{t}|x_{1},\ldots,x_{t-1};\theta)>T\quad~{}~{}\text{with}\quad~{}~{}T=\mu_{t-\tau}+\gamma\sigma_{t-\tau}$ (1)
@@ 3.3 Boundary refinement
$\mathcal{B}$: List of final boundary positions
$\mathcal{B}\leftarrow$ [i for i in range(length(tok)) if $-\log(P(\texttt{tok}[i]))>$ $T$] $\triangleright$ Boundary identification
6:for i in $\mathcal{B}$)) do
@@
8:     $\mathcal{B}[i+1]\leftarrow\arg\max_{\hat{\beta}\in(\alpha,\beta]}f(A,\{\alpha,\hat{\beta}\})$triangleright Boundary refinement
@@
While surprise-based segmentation provides an effective initial estimate of event boundaries, we make the key observation that the utility of elements within an event, during memory recall, depends on their likelihood of being utilised by the current query. Therefore, we theorise that memory recall will be most efficient with high intra-event similarity between keys while maintaining low inter-event similarity. For instance, see the similarity of groups in Fig. 2. To further ensure this, we introduce a boundary refinement step that looks to optimise this objective. Such an objective is typically optimised in the context of graph-clustering, hence we express this refinement process in a graph-theoretic manner. To achieve this, we treat the similarity matrix between all keys of an attention head $h$ within the local context window for tokens $x_{1},x_{2},...,x_{n}$ as an adjacency matrix. We define the adjacency matrix $A^{h}$ as
@@
$f_{M}(A^{h},\mathcal{B})=\frac{1}{4m}\sum_{i,j}\left[A^{h}_{ij}-\frac{\sum_{i}A^{h}_{ij}\cdot\sum_{j}A^{h}_{ij}}{2m}\right]\delta(c_{i},c_{j})$ (3)
@@
$f_{C}(A^{h},\mathcal{B})=\min_{S\in V}\frac{\sum_{i\in S,j\notin S}A^{h}_{ij}}{\min(\text{vol}(S),\text{vol}(V\setminus S))},\;\;\;\;\text{with}\;\text{vol}(S)=\sum_{i,j\in S}A_{ij},\;\text{vol}(V\setminus S)=\sum_{i,j\notin S}A_{ij}$ (4)
where $S=\{b_{i},b_{i}+1,...,b_{i+1}\}$ is a subset of all nodes $V=\{b_{1},b_{1}+1,...,b_{k}\}$ in the induced graph, with $b_{i}\in\mathcal{B}$. Lower conductance values indicate better community structure. Our boundary refinement algorithm iteratively adjusts the initial surprise-based boundaries to optimise these metric functions. While our best results are achieved using modularity, we also include comparisons with conductance-based boundary refinement to provide a comprehensive analysis. The overall process is summarized in Algorithm 1.
This algorithm first identifies initial boundaries based on the surprise threshold $T$, then refines these boundaries by finding the optimal position $\hat{\beta}$ between each pair of consecutive initial boundaries $(\alpha,\beta)$ that optimises the chosen metric function $f$ (either maximising modularity or minimising conductance). This process ensures that the final segmentation (1) captures points of high surprise and (2) optimises for coherent information grouping. The boundary identification step incurs negligible computational cost, as it only evaluates existing LLM outputs. The time complexity of Algorithm 1 is dominated by the boundary refinement step, which has an overall complexity of $\mathcal{O}(kn)$, where $k$ is the number of initial boundaries and $n$ is the sequence length (for details see Appendix C.1).
@@ 3.4 Memory Retrieval
The second stage of our retrieval process introduces another buffer, which we refer to as the contiguity buffer, designed to maintain temporal context. Implemented as a queue of size $k_{c}$, this buffer promotes temporal relationships in retrieval. When an event is retrieved, we also enqueue its neighboring events (within ±$\pm n$ positions in the original sequence) into this buffer. This mechanism enables the LLM’s ‘‘induction’’ attention heads to exhibit the contiguity and asymmetry effects discussed in Section 2.2. The queue structure allows for a natural decay of temporal context as new events are processed, with older or repeated events being dequeued as new ones are added. In total, $k=k_{s}+k_{c}$ events are added to the context window, striking a balance between relevance and temporal relationships in a manner analogous to human episodic memory retrieval. Note that each layer retrieves and attends to these $k$ events individually, allowing it to potentially focus on different parts of the context.
@@ 4.1 Performance of EM-LLM on long-context tasks
Base LLM Method LongBench $\infty$-∞ -Bench
@@
Table 1: EM-LLM performance on LongBench (grouped tasks) and $\infty$-Bench compared to our baseline InfLLM. S: surprise threshold, SM: surprise threshold and refinement with modularity, S+C: surprise threshold and contiguity buffer, SM+C: surprise, refinement and contiguity buffer. Each row indicates the number of local + retrieved tokens (eg. 4k+2k) used for both InfLLM and EM-LLM.
Results on both benchmarks (Table 1) show that our method is able to improve on InfLLM across 5 different base LLMs, $80\%$ of individual task groups of LongBench and on the overall average. Note that the table shows the best single method in terms of overall performance for each ablation. Looking at individual task performance across all experiments, EM-LLM is able to surpass InfLLM in all tasks (see Appendix A.1). Notably, we see an especially large jump in performance in the retrieval (Passage, KV, Passkey, Number) and QA (Narrative, Qasper, MultiField, Hotpot, 2Wiki and Musique) tasks across all ablations, with up to a $40\%$ and $29.7\%$ improvement over InfLLM respectively. Such tasks require the model to identify and retrieve specific information within the input sequence, a challenging test for the model’s ability to accurately recall a wide range of detailed information from a large context concurrently. This substantial improvement highlights the effectiveness of our event segmentation method in enhancing long-term memory recall and retrieval accuracy in LLMs.
@@
To evaluate EM-LLM against prominent methods for handling long contexts, we compared its performance on LLaMA 3.1-8B with two different RAG approaches, including the current SOTA NV-Embed-v2 retriever [Lee et al., 2024], as well as with the brute-force baseline of processing all tokens directly within the LLM’s softmax attention (full-context). Across most tasks in our benchmarks, EM-LLM outperformed both RAG and full-context methods (Fig. 1 and Appendix A.2), exceeding the performance of NV-Embed-v2 by $30.5\%$ on LongBench and by $11.5\%$ on $\infty$-Bench.
This significant performance boost over RAG can be attributed to EM-LLM’s ability to retrieve and incorporate relevant information at each layer individually, rather than relying on a single retrieval step as in RAG (for an illustration, see Supp. Fig. 5). By accessing more specific and contextually relevant information through layer-wise key-value retrieval, EM-LLM effectively addresses RAG’s limitations in precision and lower overall performance [Li et al., 2024b]. Additionally, EM-LLM’s hierarchical attention avoids the issue of diluted attention in large context windows that affects full-context models, enabling it to outperform both RAG and full-context LLMs on the LongBench dataset. Furthermore, EM-LLM demonstrated remarkable scalability by achieving $100\%$ accuracy on the Passkey.Retrieval task with sequences up to $5$,$120$k tokens, far beyond the practical limits of full-context LLMs. This highlights EM-LLM’s efficiency in handling extremely long contexts, positioning it as a powerful alternative for long-context processing.

### v3

# Human-inspired Episodic Memory for Infinite Context LLMs

@@ Abstract
Large language models (LLMs) have shown remarkable capabilities, but still struggle with processing extensive contexts, limiting their ability to maintain coherence and accuracy over long sequences. In contrast, the human brain excels at organising and retrieving episodic experiences across vast temporal scales, spanning a lifetime. In this work, we introduce EM-LLM, a novel approach that integrates key aspects of human episodic memory and event cognition into LLMs with no fine-tuning, enabling them to handle practically infinite context lengths while maintaining computational efficiency. EM-LLM organises sequences of tokens into coherent episodic events using a combination of Bayesian surprise and graph-theoretic boundary refinement in an online fashion. When needed, these events are retrieved through a two-stage memory process, combining similarity-based and temporally contiguous retrieval for efficient, human-inspired access to relevant information. Experiments on the LongBench and $\infty$-Bench benchmarks demonstrate EM-LLM’s superior performance, consistently outperforming the state-of-the-art retrieval model InfLLM across various baseline LLMs. In addition, EM-LLM outperforms its popular counterpart, RAG, in a wide range of tasks, while requiring similar resources. Notably, EM-LLM’s performance even surpasses full-context models in most tasks, while successfully performing retrieval across 10 million tokens -- a scale computationally infeasible for such models. Finally, our analysis reveals strong correlations between EM-LLM’s event segmentation and human-perceived events, suggesting parallels between this artificial system and its biological counterpart, thereby offering a novel computational framework for exploring human memory mechanisms.
@@ 1 Introduction
Figure 1: Top: EM-LLMS (surprise only) vs. RAG (NV-Embed-v2 retriever) vs. full-context, with LLaMA-3.1-8B as the base LLM, evaluated on LongBench. Bottom: Comparison of various long-sequence methods (sorted based on their context window length) on an extended version of $\infty$-Bench’s Retrieve.PassKey. Baseline data taken from Ding et al. [2024].
@@
We propose EM-LLM, a novel architecture integrating crucial aspects of event cognition and episodic memory into Transformer-based LLMs through three key innovations (a, b and c). For memory formation, we segment input token sequences into memory units representing episodic events. The boundaries of these units are (a) initially determined using the model’s surprise level during inference, then (b) refined to maximize within-unit cohesion and cross-unit separation (see Section 3.2). This refinement leverages graph-theoretic metrics, treating attention key similarity as a weighted adjacency matrix, and aims to enhance efficient information recall in complex, long-context tasks: by consolidating related information into single units, we seek to minimize the number of units needed for event-specific recall. The resulting memory formation process is computationally efficient: surprise-based segmentation requires no additional computation, and refinement complexity is $\mathcal{O}(nm)$, where $m$ is typically negligible compared to the token count $n$ in long-context tasks. For memory recall, (c) our approach combines similarity-based retrieval with temporal contiguity and asymmetry mechanisms, building on recently discovered parallels between LLMs and human sequential information retrieval patterns [Ji-An et al., 2024]. This method therefore ensures efficient information access while replicating temporal dynamics from human free recall studies [Howard and Kahana, 2002], and enhancing performance on tasks requiring temporal reasoning. See Appendix E.2 for analysis of EM-LLM’s architectural contributions.
Performance:
@@
We show that our method is scalable and significantly outperforms the SOTA retrieval model InfLLM [Xiao et al., 2024a], as well as RAG and full-context methods, on the widely-used LongBench [Bai et al., 2023] and $\infty$-Bench [Zhang et al., 2024] benchmarks designed for long-context tasks (see Fig. 1). Furthermore, we perform successful passkey retrieval across 1010M tokens, a length which is computationally infeasible for current full-context models. To further prove our hypotheses, we then employ a series of human-annotated podcast scripts to show that information in LLM attention heads can be semantically grouped in a way that correlates with the event structure perceived by humans. Therefore, LLM-perceived surprise can indeed serve as a proxy for the cognitive signals that drive human event segmentation, as confirmed by previous studies [Kumar et al., 2023]. Finally, using the long-context PG-19 dataset [Rae et al., 2020], which comprises a diverse corpus of English books, we evaluate the effectiveness of our segmentation method for grouping relevant information and assess the performance of different boundary refinement objectives.
@@ 2.1 Long-context in LLMs
Recently, several approaches have been proposed to extend the context window of Transformer-based models. These include methods that address the limited representational capacity of softmax attention, and its quadratic computational and memory cost [Katharopoulos et al., 2020, Munkhdalai et al., 2024]. Other methods target the poor extrapolation of typical positional encodings to out-of-distribution context lengths [Kazemnejad et al., 2024]. The latter is evident in most widely used methods, including the original absolute positional encodings [Vaswani et al., 2017] and the more recent relative positional encodings, such as the Rotary Positional Embeddings (RoPE) [Su et al., 2024]. To address this, some methods propose scaling of the rotation angles [Chen et al., 2024a] or the base constant in RoPE [Xiong et al., 2023, Liu et al., 2024b, Peng et al., 2024, Ding et al., 2024]. Others, scale positions without affecting the embedding function [Press et al., 2021, Chen et al., 2023, Jin et al., 2024], explore alternative strategies such as KERPLE [Chi et al., 2022] and FIRE [Li et al., 2024a] or adopt relative position mechanisms from certain LMs like T5 [Raffel et al., 2020].
@@
Concerning computational efficiency and diluted attention, successful approaches propose methods for general improvements to Transformer efficiency through optimised computations [Dao, 2024, Han et al., 2024a, Aminabadi et al., 2022, Kwon et al., 2023, Liu et al., 2024c, Brandon et al., 2023] or compression techniques [Nawrot et al., 2024, Zhang et al., 2023], as well as training methods tailored for long-context scenarios [Zhu et al., 2024, Chen et al., 2024b]. Another direction is the utilisation of retrieval-based methods, the vast majority of which relies on a vector database that keeps a key-value cache and scalable approximations of k-nearest neighbors (k-NNs) to perform lookups [Wu et al., 2022, Tworkowski et al., 2023, Bertsch et al., 2023]. Interestingly, since using a key-value cache with k-NN lookup can be seen as an approximation of applying softmax attention to the full token sequence (see Appendix F.1), k-NN retrieval methods can be used without fine-tuning [Bertsch et al., 2023]. For an exception that does not rely on k-NNs, see Wang et al. [2023].
@@
A recent and interesting variant of k-NN retrieval involves retrieving large groups of tokens, rather than individual ones. Models that rely on this approach include SLED [Ivgi et al., 2023] and the more recent InfLLM [Xiao et al., 2024a], which achieves SOTA performance on long-context benchmarks. InfLLM segments the entire context length into fixed-size memory units and employs k-NN lookup using the tokens with the highest accumulated scores per unit. The latter can be seen as a form of hierarchical attention in models that use such retrieval, as illustrated in Fig. 2. While group-based retrieval represents a promising direction, our approach significantly advances this concept by dynamically determining token groupings in a manner akin to human memory formation. This effectively addresses a fundamental limitation of InfLLM’s fixed-size segmentation and enables more adaptive and context-sensitive processing of extended information.
@@ 3.2 Memory formation via Surprise
$-\log P(x_{t}|x_{1},\ldots,x_{t-1};\theta)>T\quad~~\text{with}\quad~~T=\mu_{t-\tau:t}+\gamma\sigma_{t-\tau:t}$ (1)
@@ 3.3 Boundary refinement
[…]
6:for i in range(length($\mathcal{B}$)) do
@@
[…]
@@
While surprise-based segmentation provides an effective initial estimate of event boundaries, we make the key observation that the utility of elements within an event, during memory recall, depends on their likelihood of being utilised by the current query. Therefore, we theorise that memory recall will be most efficient with high intra-event similarity between keys while maintaining low inter-event similarity. For instance, see the similarity of groups in Fig. 2. To further ensure this, we introduce a boundary refinement step that looks to optimise this objective. Such an objective is typically optimised in the context of graph-clustering, hence we express this refinement process in a graph-theoretic manner. To achieve this, we treat the similarity matrix between all keys of an attention head $h$ within the local context window for tokens $x_{1},x_{2},...,x_{n}$ as an adjacency matrix $A^{h}$:
@@
$f_{M}(A^{h},\mathcal{B})=\frac{1}{4m}\sum\nolimits_{i,j}\left[A^{h}_{ij}-\frac{1}{2m}{\sum\nolimits_{i}A^{h}_{ij}\cdot\sum\nolimits_{j}A^{h}_{ij}}\right]\delta(c_{i},c_{j})$ (3)
@@
$f_{C}(A^{h},\mathcal{B})=\min_{S\in V}\frac{\sum_{i\in S,j\notin S}A^{h}_{ij}}{\min(\text{vo}(S),\text{vo}(V\setminus S))},\;\;\text{with}\;\text{vo}(S)=\sum_{i,j\in S}A_{ij},\;\text{vo}(V\setminus S)=\sum_{i,j\notin S}A_{ij}$ (4)
where $S=\{b_{i},b_{i}+1,...,b_{i+1}\}$ is a subset of all nodes $V=\{b_{1},b_{1}+1,...,b_{k}\}$ in the induced graph, with $b_{i}\in\mathcal{B}$. Lower conductance values indicate better community structure. Our boundary refinement algorithm sequentially adjusts the initial surprise-based boundaries to optimise these metric functions. While our best results are achieved using modularity, we also include comparisons with conductance-based boundary refinement to provide a comprehensive analysis. The overall process is summarized in Algorithm 1 and further discussed in Appendix E.3.
This algorithm first identifies initial boundaries based on the surprise threshold $T$, then refines these boundaries by finding the optimal position $\hat{\beta}$ between each pair of consecutive initial boundaries $(\alpha,\beta)$ that optimises the chosen metric function $f$ (either maximising modularity or minimising conductance). This process ensures that the final segmentation (1) captures points of high surprise and (2) optimises for coherent information grouping. The boundary identification step incurs negligible computational cost, as it only evaluates existing LLM outputs. The time complexity of Algorithm 1 has an overall complexity of $\mathcal{O}(nm)$, where $n$ is the $n$ is the sequence length and $m$ is the chunk size selected to process the sequence (for details see Appendix C.1).
@@ 3.4 Memory Retrieval
The second stage of our retrieval process introduces another buffer, which we refer to as the contiguity buffer, designed to maintain temporal context. Implemented as a queue of size $k_{c}$, this buffer promotes temporal relationships in retrieval. When an event is retrieved, we also enqueue its neighboring events (within $\pm n$ positions in the original sequence) into this buffer. This mechanism enables the LLM’s ‘‘induction’’ attention heads to exhibit the contiguity and asymmetry effects discussed in Section 2.2. The queue structure allows for a natural decay of temporal context as new events are processed, with older or repeated events being dequeued as new ones are added. In total, $k=k_{s}+k_{c}$ events are added to the context window, striking a balance between relevance and temporal relationships in a manner analogous to human episodic memory retrieval. Note that each layer retrieves and attends to these $k$ events individually, allowing it to potentially focus on different parts of the context.
@@ 4.1 Performance of EM-LLM on long-context tasks
Base LLM Method LongBench $\infty$-Bench
@@
Table 1: EM-LLM performance on LongBench (grouped tasks) and $\infty$-Bench compared to our baseline InfLLM. S: surprise threshold, SM: surprise threshold and refinement with modularity, S+C: surprise threshold and contiguity buffer, SM+C: surprise, refinement and contiguity buffer. Each row indicates the number of local + retrieved tokens (eg. 4k+2k) used for both InfLLM and EM-LLM. See Appendix D.1 for parameter choices and Appendix A.1 for more results and significance testing.
Results on both benchmarks (Table 1) show that our method is able to improve on InfLLM across 5 different base LLMs, $80\%$ of individual task groups of LongBench and on the overall average. Note that the table shows the best single method in terms of overall performance for each ablation (see Appendix A.1 for all ablations in methods). Looking at individual task performance across all ablations in methods, EM-LLM is able to surpass InfLLM in all tasks. Notably, we see an especially large jump in performance in the retrieval (Passage, KV, Passkey, Number) and QA (Narrative, Qasper, MultiField, Hotpot, 2Wiki and Musique) tasks across all ablations, with up to a $40\%$ and $29.7\%$ improvement over InfLLM respectively. Such tasks require the model to identify and retrieve specific information within the input sequence, a challenging test for the model’s ability to accurately recall a wide range of detailed information from a large context concurrently. This substantial improvement highlights the effectiveness of our event segmentation method in enhancing long-term memory recall and retrieval accuracy in LLMs.
@@
To evaluate EM-LLM against prominent methods for handling long contexts, we compared its performance on LLaMA 3.1-8B with two different RAG approaches, including the current SOTA NV-Embed-v2 retriever [Lee et al., 2024], as well as with the brute-force baseline of processing all tokens directly within the LLM’s softmax attention (full-context). Across most tasks in our benchmarks, EM-LLM outperformed both RAG and full-context methods, as well as a custom surprise-based RAG method (Fig. 1 and Appendix A.2), exceeding the performance of NV-Embed-v2 by $30.5\%$ on LongBench and by $11.5\%$ on $\infty$-Bench.
This significant performance boost over RAG can be attributed to EM-LLM’s ability to retrieve and incorporate relevant information at each layer individually, rather than relying on a single retrieval step as in RAG (for an illustration, see Supp. Fig. 5). By accessing more specific and contextually relevant information through layer-wise key-value retrieval, EM-LLM effectively addresses RAG’s limitations in precision and lower overall performance [Li et al., 2024b]. Additionally, EM-LLM’s hierarchical attention avoids the issue of diluted attention in large context windows that affects full-context models, enabling it to outperform both RAG and full-context LLMs on the LongBench dataset. Furthermore, EM-LLM demonstrated remarkable scalability by achieving $100\%$ accuracy on the Passkey.Retrieval task with sequences up to 10.210.2M tokens, far beyond the practical limits of full-context LLMs. This highlights EM-LLM’s efficiency in handling extremely long contexts, positioning it as a powerful alternative for long-context processing.